OpenStudy (cashgirl1369):

ΔABC is a right triangle in which angle b is a right angle, AB = 1, AC = 2, and BC =square root 3 . cos C × sin A =

OpenStudy (luffingsails):

Do you know that cos(angle) = opposite side/hypotenuse sin(angle) = adjacent side/hypotenuse

OpenStudy (cashgirl1369):

a little

OpenStudy (luffingsails):



OpenStudy (cashgirl1369):

i lost you

OpenStudy (luffingsails):

Do you understand the figure I put up?

OpenStudy (cashgirl1369):

a lil

OpenStudy (luffingsails):

Focus on the angle at C. the cosine of that angle is defined as the opposite side divided by the hypotenuse. We know both of those lengths. So, cos(C) = opposite side/hypotenuse cos(C) = 1/2

OpenStudy (luffingsails):

Now angle at A. The sine of that angle is defined as adjacent side divided by hypotenuse. We also know both of those lengths. So, sin(A) = adjacent side/hypotenuse sin(A) = 1/2 Then cos(C) * sin(A) = (1/2) * (1/2) = 1/4

OpenStudy (cashgirl1369):

i still dont understand

OpenStudy (luffingsails):

Do you not understand what adjacent side and opposite side means? What doesn't make sense?

OpenStudy (cashgirl1369):

everything i dont understand how to do any of this